OpenStudy (sasogeek):

if i add (x to one) and subtract one from (x plus root of one squared), i get a number twice as big as (x minus itself), what is x?

OpenStudy (aravindg):

First form the equation using the info given

OpenStudy (sasogeek):

go right ahead :)

OpenStudy (aravindg):

lol why not you try first :)

OpenStudy (sasogeek):

alright... this is what i get :P lol uhmm... add x to one, so, x+1 and subtract one from x plus root of one squared, (x+1)+(x+1^2-1) that's equal to a number twice as big as x minus itself, so (x+1)+(x+1^2-1)=2*(x-x) correct?

Parth (parthkohli):

(Hey, congrads!)

OpenStudy (aravindg):

yes

OpenStudy (sasogeek):

nope! LOOOL, look at this line :P and subtract one from x plus root of one squared, (x+1)+(x+1^2-1) where's the root of one squared? it's only one squared, there's no root :P

OpenStudy (aravindg):

\[\sqrt{1^2}=1\]

OpenStudy (aravindg):

No difference :)

Parth (parthkohli):

The question is kinda confusing. What does it exactly say? What are you doing with \(x + 1\) and \((x + \sqrt{1})^2 - 1\)?

OpenStudy (sasogeek):

that makes all the difference in the initial equation. i think it matters though result may be the same, lol

OpenStudy (sasogeek):

@ParthKohli the full equation is what i have written down, (x+1)+(x+1^2-1)=2*(x-x) but then this part is supposed to be \(\large (x+ \sqrt{1^2}-1) \) :)

Parth (parthkohli):

Is the question this?\[(x + 1) + (x + \sqrt{1})^2 - 1 = 2(x - x)\]If so, you have \((x + 1) + (x + \sqrt{1})^2 - 1 = 0\)

OpenStudy (sasogeek):

tbh though i shouldn't have let that one out lol, it was supposed to be left for you to figure out :P

OpenStudy (aravindg):

actually When you frame the equation it should be important \[(x+1)+(x+\sqrt{1}^2-1)=2 * (x-x)\]

Parth (parthkohli):

Hmm, but I don't think that it'd be \(\sqrt{1}^2\), but if you insist so...

OpenStudy (sasogeek):

root of 1 squared has only 1 meaning lol, \(\sqrt{1^2} \)

Parth (parthkohli):

No... this can't be

Parth (parthkohli):

You will have two solutions then, I guess.

OpenStudy (sasogeek):

nope, just one

OpenStudy (aravindg):

just one

OpenStudy (aravindg):

\[\sqrt{1^2}=1\]

OpenStudy (aravindg):

Okay so why not solve it straight away @sasogeek ?

Parth (parthkohli):

Since \(\sqrt{1^2} = \sqrt{1} = \pm 1\) you may have to solve both equations\[(x + 1) +(x - 1)-1 = 0\]\[(x+1) + (x + 1) -1 =0\]

OpenStudy (aravindg):

omg \[\sqrt{1} \neq \pm 1\]

Parth (parthkohli):

@AravindG Why not?

OpenStudy (sasogeek):

\(\large \sqrt{1^2}=1^{2*\frac{1}{2}}=1 \)

OpenStudy (aravindg):

Because square root is a function having range as positive real numbers

Parth (parthkohli):

\[(-1)^2 = 1 \Rightarrow -1 = \sqrt{1}\]\[(1)^2 = 1 \Rightarrow 1 = \sqrt{1}\]\[\therefore \sqrt{1} = \pm 1\]@AravindG The domain is positive real number, not the range

Parth (parthkohli):

Or actually, the domain is nonnegative real numbers.

OpenStudy (dumbsearch2):

Wow.

OpenStudy (aravindg):

@ParthKohli you are having a serious misconception on square roots .

OpenStudy (sasogeek):

you don't have to spam

Parth (parthkohli):

@AravindG Wasn't my demonstration enough?

OpenStudy (sasogeek):

can we just solve my problem and take this to another thread, lol

OpenStudy (anonymous):

so where was this problem from?

OpenStudy (aravindg):

I can explain it to you separately not here

OpenStudy (sasogeek):

my head, @Peter14

OpenStudy (aravindg):

Lets finish this problem first

OpenStudy (anonymous):

you have an interesting head.

OpenStudy (sasogeek):

lol why thank you :)

OpenStudy (aravindg):

ok @sasogeek proceed :)

Parth (parthkohli):

If you want to neglect one root of \(1\), you may. Just consider \(1\) as the root of 1 if you want to. But seriously, this equation has two possible solutions.

OpenStudy (dumbsearch2):

LOL @Peter14

OpenStudy (aravindg):

@ParthKohli I can make a separate thread for your query .

OpenStudy (sasogeek):

it's not that hard u guys -_-, it boils down to 2x+1=0, hard? i don't think so xD

OpenStudy (sasogeek):

x=\(-\frac{1}{2} \)